• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Voutier

    Log entry time 16:08:44 on October15,2004

    Entry number 134307

    keyword=Shift Summary


    New Crew on shift: Amarian, Zhu, Voutier
    -----------------

    8:00 Shift change while beam has been taken away.

    8:04 Beam restore to the hall, MCC wants to do OTR testat 1C12

    8:09 Start production (MCC will call later for OTR)

    8:50 Acquisition problems (kcoda twice and reboot all ROCs)

    9:09 Back to data production

    10:28 Back to acquisition troubles

    11:03 DVCS production start again

    11:08 Beam problems / a crate in the accelerator

    11:34 Pedestal run in DVCSPed configuration

    11:54 Back to data production

    12:25 Notice that the Compton rate is zero.

    12:54 Stop run, any new run is stuck (ROC18 problem)

    13:23 Succeed to start a run / Alexandre is on his way to cure the
    problem

    14:15 Proton array timinig changes by 24ns (entry 134300)

    14:18 Back to production run

    Run History
    -----------

    #3715 DVCS production, good run, 200 k
    #3716 - #3720 Junk, coda problems
    #3721 DVCS production, good run, 200 k
    #3722 DVCS production, good run, 200 k
    #3723 - #3727 Junk, coda problems
    #3728 DVCS production, good but short run stop, 20 k
    #3729 Pedestal run
    #3730 DVCS production run, good run, 200 k
    #3731 - #3736 Junk, acquisition troubles
    #3737 DVCS production, good run, 200 k
    #3738 Junk
    #3739 DVCS production, PA TIMING CHANGES, good run, 200 k
    #3740 DVCS production, good run, 200 k
    #3741 DVSC production ... goes to next shift